# Personalized Generative Low-light Image Denoising and Enhancement

### v1

@@ Appendix E More Implementation Details
We compared the restored images generated with the encoder-based method with the proposed physical buffer-guided method on our real-captured photo set. The quantitative and visual comparisons are presented in Table 9 and Fig. 15, respectively. The results demonstrate that guiding the diffusion model using ID physical buffers produces significantly better identity preservation than the encoder-based approach. One major reason for this is that the ID physical buffers serve as robust conditions, enabling the extraction of comprehensive facial features from the gallery. In contrast, the encoder-based method is less focused on facial features and is often influenced by irrelevant factors such as background and hairstyle. Consequently, the latent codes extracted by the encoder are less effective in guiding the diffusion model to the personalized generative space compared to the ID physical buffers.
Appendix G Gallery Photos
Fig. 19 presents the sampled gallery photos for each test subject included in the main paper and supplementary material. Additionally, the gallery photos of the real-captured male subject are displayed in Fig. 9.
Figure 12: Visual comparison on CelebAMask-HQ test set under extreme conditions (10 PPP).
Figure 13: Visual comparison on CelebAMask-HQ test set under extreme conditions (5 PPP).
Figure 14: Visual comparison on the real-world captured SUBJECT under extreme conditions (zoomed-in area in the input is scaled for visual purposes).
@@ Appendix G Gallery Photos
Figure 16: Visual comparison on CelebAMask-HQ test set. Zoom in for the best visual experience.
Figure 17: Visual comparison on CelebAMask-HQ test set. Zoom in for the best visual experience.
Figure 18: Visual comparison on CelebAMask-HQ test set. Zoom in for the best visual experience.
Figure 19: Gallery samples of testing persons appeared in the main paper and supplementary materials.

### v2

@@ Appendix C More Implementation Details
We compared the restored images generated with the encoder-based method with the proposed physical buffer-guided method on our real-captured photo set. This is a special semi-real set, in which we take photos of some printed paper photos at night. We prepare this semi-real set because The ground-truth will be available to be compared with. The quantitative and visual comparisons are presented in Table 6 and Fig. 11, respectively. The results demonstrate that guiding the diffusion model using ID physical buffers produces significantly better identity preservation than the encoder-based approach. One major reason for this is that the ID physical buffers serve as robust conditions, enabling the extraction of comprehensive facial features from the gallery. In contrast, the encoder-based method is less focused on facial features and is often influenced by irrelevant factors such as background and hairstyle. Consequently, the latent codes extracted by the encoder are less effective in guiding the diffusion model to the personalized generative space compared to the ID physical buffers.
Appendix E Gallery Photos
Fig. 13 presents the the gallery photos for the test subjects included in the Fig. 6 of the main paper.
@@ Appendix E Gallery Photos
Figure 12: Additional real-captured visual comparison.
Figure 13: Gallery photos.
Figure 14: Additional visual comparison.
Figure 15: Additional visual comparison.